OpenStudy (anonymous):

What is the simplified form of 24 y to the fifth power over 15 x to the eighth power divided by 8 y¨Csquared over 4 x to the fourth power ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{ 24 }{ 15 } and \frac{ 4 }{ 8 }\] can simplify to \[\frac{ 8 }{ 5 } and \frac{ 1 }{ 2 }\] \[\frac{ 8y ^{2} }{ 5x ^{8} } \times \frac{ 4x ^{4} }{ 8y ^{2} }\] Both 8 y squared's cancel out, so just multiply and get \[\frac{ 4x ^{4} }{ 5x ^{8} }\] Division in fractions with powers is simply subtraction, so \[\frac{ 4 }{ 5x ^{4} }\]